I need some help with this...
A balanced reaction equation must have the same nu
Answer #1

The given equation is as

C2H5OH + 3O2 --> 2CO2 + 3H2O

No. of atoms of each elements are as

Reactants Products
No. of carbon atoms = 2 No. of carbon atoms = 2
No. of hydrogen atoms = 5 + 1 = 6 No. of hydrogen atoms = 2 x 3 = 6
No. of oxygen atoms = 3 x 2 + 1 = 7 No. of oxygen atoms = 2 x 2 + 3 = 7

In both sides of the reaction, no. of atoms of each elements are same. Therefore, the given reaction is balanced.

Is the reaction balanced? -Yes